OpenStudy (girlgoyleh.):

WILL GIVE MEDALS! The width of a rectangle is 6 in. less than its length. The perimeter is 68 in. What is the width of the rectangle?

OpenStudy (anonymous):

let the length = l in then the width w=l-6 in then use the perimeter rule for a rectangle

OpenStudy (girlgoyleh.):

i got 14, but then i did it again and got 11!

OpenStudy (girlgoyleh.):

1-6=-5

OpenStudy (anonymous):

let x= length, y= width then, y=x-6 eq1 and for the perimeter we have 2x+2y=68 eq2 Two equations and two unknowns, perfectly solvable system. My suggestion would be to substitute equation 1 into 2 and solve for x, then use equation 1 again to solve for y

OpenStudy (girlgoyleh.):

it is 14, i got it right! Thanks both of you :)